Problem 3-11

A manager of a store that sells and installs spas wants to prepare a forecast for January, February, and March of next year. Her forecasts are a combination of trend and seasonality. She uses the following equation to estimate the trend component of monthly demand: Ft = 90 + 2t, where t = 0 in June of last year. Seasonal relatives are 1.08 for January, .87 for February, and 1.01 for March. What demands should she predict? (Round your answers to 2 decimal places.)

Month Forecast
January of the next year
February of the next year
March of the next year
Answer #1

in June Last year t=0

July,last year t=1

January, this year t=7

January, next year t-19

Feb,next year t=20

Mar, next year t=21

hence forecast for Jan next year=Sesonality component*(90+2*t)

=1.08*(90+2*19)=138.24

Feb Next year=0.87*(90+2*20)= 113.1

March,next year= 1.01*(90+2*21)=133.32

.

.

I have given complete and detailed solution to your problem, in case of any query you can comment and

PLEASE DO GIVE A THUMBS UP